p,p+7,p+14,.... is arithmetic sequence

Find n term and Find the smallest value of p for which 2021 is a term in the sequence.


1
Expert's answer
2022-03-19T02:35:35-0400

"p,p+7,p+14,..."  is arithmetic sequence

"a_1=p, a_2=p+7, a_3=p+14, d=a_2-a_1=7\\
a_n=2021\\
a_n=a_1+d(n-1)\\
2021=p+7(n-1)\\
2021=p+7n-7\\
7n=2028-p\\
n=\frac{2028-p}{7}\\ n\in N\\
n=\frac{2023+5-p}{7}\\
n=289+\frac{5-p}{7}\\
5-p=7k, k\in Z\\
p=5-7k\\
k=0, p=5\\n=289\\
a_{289}=2021, p=5"